OpenStudy (anonymous):

could someone expand this expression? y= 5x-2.. substitute in the equation 2x^2 + y^2 - 4x - 14y + 48..

OpenStudy (nirmalnema):

\[2x^2+(5x-2)^2-4x-14(5x-2)+48\]

OpenStudy (nirmalnema):

\[2x^2+(25x^2-20x+4)-4x-70x+28+48\]

OpenStudy (nirmalnema):

\[27x^2-94x+80\]

OpenStudy (nirmalnema):

@silverxx do u get it

OpenStudy (anonymous):

its 2 and 40/27

OpenStudy (anonymous):

wait.. you have any idea about tangent lines??

OpenStudy (nirmalnema):

yup ...

OpenStudy (anonymous):

oh please help./.

OpenStudy (nirmalnema):

sure......i will try,,

OpenStudy (anonymous):

2x^2 + y^2 - 4x - 14y + 48 = 0 is perpendicular tp 5x - 2y + 4 = 0

OpenStudy (anonymous):

what i did first was getting the dy/dx of the first equation then that serves as the m.

OpenStudy (anonymous):

then the second equation convert to the expression y= mx + b

OpenStudy (anonymous):

then equate both slope :)

OpenStudy (nirmalnema):

what do u want to do in the que??

OpenStudy (anonymous):

and then im lost

OpenStudy (anonymous):

derive to tese answers 2x = 5y - 28 = 0 and 15x - 6y + 22 = 0

OpenStudy (anonymous):

and 2x + 5y - 46 = 0 and 15x - 6y + 32 = 0

OpenStudy (nirmalnema):

can you pls tell me the whole que....i m little bit confused

OpenStudy (anonymous):

find the equations of the tangents and normals.. 2x^2 + y^2 - 4x - 14y + 48 = 0 is perpendicular to 5x - 2y + 4 = 0

OpenStudy (nirmalnema):

given eq is of a circle or what??

OpenStudy (anonymous):

it wasnt mentioned..

OpenStudy (anonymous):

curve i think

OpenStudy (nirmalnema):

ok...

OpenStudy (nirmalnema):

sorry i m not getting the ans,,

OpenStudy (anonymous):

its okay but wait hehe

OpenStudy (anonymous):

how can i square y = 5x+2?

OpenStudy (nirmalnema):

y^2=(5x+2)^2 y^2=25x^2+20x+4)

OpenStudy (nirmalnema):

by squaring both the side.

OpenStudy (anonymous):

the dy/dx of 2x^2 + y^2 - 4x - 14y + 48?

OpenStudy (nirmalnema):

what do u get?

OpenStudy (anonymous):

- 4x + 4/ 2y- 14

OpenStudy (nirmalnema):

yup i also get this..

OpenStudy (anonymous):

can you equate that to 2/5?

OpenStudy (anonymous):

just to make sure i have same answer with yours